Shift report 20011228 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Dec 28 2001 - 23:59:39 EST

  • Next message: Apache: "Shift report 20011229 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Dec 28 23:59:38 2001, Shift: 16:00-24:00
    Supervisor: fv, with: db (~16-18)
    Log book#11 pages 106 - 111
    
    Shift Summary:
    ------------------------------------------------------------
    0) Run summary:
    
    	6187  9920 MRS 36300 FS  2 hours.
    	6188  4979 MRS 18336 FS  1 hour
    	6189  still going ( 5182 MRS at 23:45 19888 FS)
    
            Beam data hours 4
    
    	A pretty good store once it got going. 
    
    o) Accomplishment
    Got some work done for trigger inel counters
    Add tdc's to data strream
    Added histograms to supermon
    
    Problem:Run 6186 stooped at beginning of run.
    
    16:15 Went to pp run planning meeting leaving Dana to mend another
    1.5 hour.
    
    18:00 Back from meeting - 
         The l1458A had turned off. Noticed by DB who had
    	turned back the Drift voltages etc..
    	Turned back Si-det!
    	Turned T3,T4,T5 and RICH back on badtrip which was started.
    	Rhic preparing for a physics ramp.
    
    Plan for shift:
    
    a) take physics data.  (x)
    b) Wire in the Inel timing counters (all 32 channels) to TDC's
       Dana had already strung the cables (i.e. most of the work)
    c) Add these channels to the DAQ readout. (x)
    d) Update the defaultmap.txt setting (X)
    e) Update BrRawInputData for this enhanced structure. (x)
    f) Take another look at RHIC clock timing and the bunch crossing, in order
       to prepare for the bunch polarizations setup.-- will wait for tomorrow ..
    g) don't forget the dinner in the fridge. (x)
    
    18:40 MCR called to tell they have large losses around 2 o'clock
          outgoing triplet. This may be cause for loss of HV mainframe.
          They are attepmting to steer around pp2pp ?? (roman pots claims to be out)
    
    
    19:00 Managed to get a ramp (Y/B ions good 16 and 13e11) but no collisions yet.
    
    19:20 Good store (intensity as good as it gets these early days.)
    
          Started run 6187 
          Y/B  16/13 e11 ions
          Inel rate 2K
          Trigger 6 rate 6/sec.
    
         Typical currents in Drift cahmbers
         T3:  1.5 - 1.5  2.2 microAmp
         T4:  0.86 - 0.60 - 0.7
         T5-A1  0.90,0.8 , 0.5
         T5-A3  1.34,0.3,0.6
    
         much lower than in Au-Au running (of course this is at <=1%lum.)
    
         A couple of spectr/hits dist in DC looks peculiar
         T3-1-U2 has large low time tail as well as extra hits in wires 32-35
         T3-1-V1 has a big bump below the 'real timing peak.
    
    19:30 MCR attempted steering, but no improvement. See fig in logbook.
          They nore a 4 sec. period on the v124 scalers for the inel counters.
          Not seen in other detectors.
          This is also seen by pp2pp.
    
    20:20 MCR announced pol measurement
          Stopped run 6187
    
    20:30 started RUN 6188.
    21:31 stopped RUN 6188.
          Increase DC anode voltages by +20 V (all) currents are still low.
          The last used HV settings are in the logbook 10 (not here away for copying)
          Eff judging from reconstruction of 6188 seems low
          T4 tracks / matched T1-T2 == 1/10. (would have though it should be ~1/3 at this
          low pt setting near the reference momentum. Calibrations taken from run 5260
          (latest avalibale from Au run)
    21:35 Start run 6189
    23:30 MCR called to ask if we wanted to dump (coll rate is 1400 vs 2000 at start
          so said we wanted to wait a little while.
    23:50 Announce will keep store until at least 00:30
    
    
    
    	The trigger summary web page seem to behave strangly. Asked for 8 deg, 427B
    and got nothing even though the runs are in the summary, and the A polarity works just
    fine. The same happens for MRS (the newest runs).
    
    Si had tripped during the injection- Are the detectors OK??
    
    The camac scaler 11 seem not to be counting (MRS trigger).
    -- This one was easy; input connected to input of logic unit.
    
    The inel tdc setup needs more discriminators. The 16 ch used to generate logic (narrow
    pulses in incompatible with the driving of a signal to the TDC channel. There is enough 
    space to add another 16 ch driver disr for the TDC. This helps- will have to check 
    to Right side  - out of NIM slots.
    
    The inel delays are quite long - so the spectra show up toward the very end.
    Add 16 nsec to start for these tdc.(right at tdc's)
    
    
    --detailed log
    
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Dec 29 2001 - 00:00:23 EST